Oxidation numbers of C and N in CNO-. What is answer to this problem

Grade:11

1 Answers

aakash sharma
6 years ago
 
I think your compound is (nh4)2 so4 
presently, 
atomic mass of compound=2*(14+1*4)+32+16*4 
=2*18+32+64 
= 36+96 =132g 
nuclear weight of nitrogen=28g 
along these lines, 
% of nitrogen in 66% =28/(66/100 *132) %=32 % approx